please help me solve these problems. I cannot relate to the studies of my son. Thank you very much!
1. My uncle is 5 yrs. older than trice my eldest brothers age. My eldest brother is 17 yrs. old. How old is my uncle?
2. Lynn has 2 1/4 meters of white cloth. She used 5/3 meters for a dress. How much cloth does Lynn have left?
3. There are 54 pupils in Grade 3 class. Five ninths of them are girls and 1/3 of the girls wear black shoes. How many girls wear black shoes?
4. If I subtract 12 from the number I am thinking of and then, multiply the difference by 5, the result is 65. What is the number I am thingking of?
5. Bea skipped counted by 4 from 10, what number did she reach when she skip counted 12 times.
6. Elmin is skip counting by 8 as he skips rope. He said 15 on her first skip. How many skips have he made when he says 111?
7. What is the perimeter of rectangle whose dimensions are 14 meters and 16 meters?
8. one side of a baseball diamod is 20 meters long. How far does a runner go in running once round it?

Answer by mathstutor494(120)

1. Uncle's Age = 5 yrs. older than trice my eldest brothers age
My eldest brother = 17 yrs
So Uncle's age = (3*17)+5
56 Years
2. Lynn has 2 1/4 meters of white cloth. She used 5/3 meters for a dress.
2 1/4 = 9/4
(9/4) - (5/3)
Multiplying both numerator & denominator of first term by 3 and second term by 4, we get ..
27/12 - 20/12
(27-20)/12
7/12
Ans : Cloth left with Lynn is 7/12 meters